A technician has been given 100ml of a 10% acetic acid solution. The pharmacist asked the technician to dilute the solution of 500 ml with sterile water and to label the solution. What percent should appear on the label?
Answer by checkley77(12844)

IF the total mix is 500ml then:
.10*100=500x
10=500x
500x=10
x=10/500
x=.02 or thre mix is 2% acid.
-------------------------------------------
IF you're adding 500ml of water then:
.10*100=(100+500)x
10=600x
600x=10
x=10/600
x=.0167 or 1.67% acid.